# Born series for s-wave scattering length and some exact results

### v1

@@ Abstract
In these notes the Born series for the $s$-wave scattering $a_{0}$ is calculated for a class of central potentials $V(r)$ up to sixth order in a dimensionless coupling strength $g$. Examples of exponentially decaying potentials as well truncated potentials involving a single length-scale $a$ are considered. In certain favorable cases the exact result for the $g$-dependent $s$-wave scattering length $a_{0}=A_{0}(g)a$ can be given in terms of special functions. The poles of $A_{0}(g)$ at increasing positive values of $g$ correspond to the thresholds, where $s$-wave bound-states occur successively.
@@ 2 Exponentially decaying potentials
In this section the Born series $A_{0}=\sum_{j=1}^{6}c_{j}g^{j}+\dots$ is computed for a variety of central potentials $V(r)$ with an exponential decay. The familiar Yukawa potential is physically well motivated in terms of an underlying one-boson exchange mechanism. The pertinent profile function $f(x)$ and ”electrostatic potential” $\Phi(x)$ are in this case
@@
With this input one derives for the Born series up to sixth order:
@@
The series for $A_{0}(g)$, with its first six terms written in eq.(8), has a finite radius of convergence $(|g|<g_{1})$, that is determined as that strength parameter $g_{1}$, where the threshold for the first $s$-wave bound-state in the Yukawa potential is reached. This particular value $g_{1}$ can be computed by solving numerically the second-order differential equation for the radial wave-function (multiplied with $x$):
@@
subject to a linear behavior $u(x)\sim x$ at small $x$. Note that $u^{\prime\prime}(x)+gx^{-1}u(x)=0$ has the regular solution in the form of a Bessel-function $u(x)=\sqrt{g\,x}J_{1}(2\sqrt{g\,x})=g\,x+{\cal O}(x^{2})$, which also displays a linear behavior. Once the critical strength $g_{1}$ is reached, the asymptotic behavior of $u(x)$ for large $x$ turn into a constant. The convergence of this process is monitored by inspecting, how the first derivative $u^{\prime}(x)$ approaches zero for large $x$, when fine-tuning more and more digits of the strength parameter $g$. As a result of this numerical exercise, one finds that the first threshold $g_{1}$ for an $s$-wave bound-state in the Yukawa potential lies at the value:
$g_{1}=1.7443138\,.$ (12)
Next, we consider a purely exponentially decreasing potential $V(r)\sim-g\,e^{-r/a}$ with the profile function
@@
Figure 1: Scattering length $A_{0}=a_{0}/a$ versus $g$ for the exponential potential $V(r)\sim-g\,e^{-r/a}$.
@@
Note that the numerical method, employed before to construct $g_{1}$ for the Yukawa potential, reproduces $g_{1}=1.4457965$ for the profile $f(x)=e^{-x}$.
As a further example, we choose an exponentially decreasing potential with a soft core:
@@
and derive the Born series up to sixth order
@@
where all expansion coefficients are in fact rational numbers. The radius of convergence $|g|<g_{1}$ of the series for $A_{0}(g)$ written in eq.(19) is determined by searching the threshold for the first $s$-wave bound-state in the potential $V(r)\sim-g\,x\,e^{-x}$. One finds with high numerical accuracy the value
$g_{1}=0.676681997\,.$ (20)
Next, we choose an exponentially decreasing potential with a very soft core, taking the profile function
@@
where all expansion coefficients are (large) rational numbers. The radius of convergence $|g|<g_{1}$ of the series for $A_{0}(g)$ in eq.(22) is again determined by searching the threshold for the first $s$-wave bound-state in the potential $V(r)\sim-g\,x^{2}e^{-x}$. One finds with high numerical accuracy the value
$g_{1}=0.216855258\,.$ (23)
Finally, one can consider an exponentially decreasing potential with a more singular behavior at small distances:
@@
The profile $f(x)=e^{-x}/x^{2}$ is too singular near the origin and therefore the coupling strength $g$ must be restricted to negative and small positive values $g<g_{\text{crit}}=1/4$. This feature concerning an instability is further elaborated on in the next section, when analyzing the truncated $1/r^{2}$-potential.
@@ 3 Truncated potentials
A well-known textbook problem [1] is the quantum mechanical scattering off a spherical potential well, represented by the profile
@@
and now the knowledge of the ”electrostatic potential” $\Phi(x)$ is needed only in the inside region $0<x<1$. With little effort, one obtains the following Born series for the $s$-wave scattering length in a spherical potential well:
@@
and, as a nice check, the Taylor expansion of the tangent-function confirms the series in eq.(27). The analytical result for $A_{0}(g)$ written in eq.(28) is shown in Fig. 2 in the region 40<g<62-40<g<62. The poles of $A_{0}(g)$ at the values $g_{j}=\pi^{2}(j-1/2)^{2}$ (where $\tan\sqrt{g_{j}}$ diverges) correspond to the well-known thresholds for the occurrence of successive $s$-wave bound-states in a spherical potential well. The first six values of this infinite sequence read
@@
Figure 2: $A_{0}$ as a function of $g$ for the spherical well potential.
It is also interesting to study the truncated $1/r^{2}$-potential with the profile function
@@
where second term in curly brackets, $\sqrt{-g}J_{\nu+1}(\sqrt{-g})/J_{\nu}(\sqrt{-g})$, means $\sqrt{g}I_{\nu+1}(\sqrt{g})/I_{\nu}(\sqrt{g})$ for $0<g<1/4$. The exact result for $A_{0}(g)$ written in eq.(33) is shown in Fig. 3 for the parameter range 5<g<1/4-5<g<1/4. As already visible in Fig. 2, and valid for any truncated potential, the scattering length approaches with increasing repulsion the hard-core value: $\lim_{g\to-\infty}A_{0}(g)=-1$. On the positive side the boundary value, before instability sets in, is given by
@@
Next, we consider the truncated Coulomb-potential with profile function
@@
Figure 3: $A_{0}$ as a function of $g$ for the truncated $1/r^{2}$-potential, restricted to the range $g\leq 1/4$.
Figure 4: $A_{0}$ as a function of $g$ for the truncated Coulomb-potential.
@@
Figure 5: $A_{0}$ as a function of $g$ for the truncated oscillator potential.
Figure 6: $A_{0}$ as a function of $g$ for the truncated linear potential.
A further interesting example is the truncated oscillator potential with the profile function
@@
As a last example we consider the truncated linear potential with the profile function
@@
4 Further solvable potential scattering problem
In passing we present and discuss another solvable scattering problem, where the potential has the shape of a squared Lorentz-curve, i.e. the profile function is
@@ 4 Further solvable potential scattering problem
$A_{0}(g)$ $\!\!=\!\!$ $g\,{\pi\over 4}+g^{2}\,{\pi\over 16}+g^{3}\,{\pi(\pi^{2}-6)\over 192}+g^{4}\,{\pi(15-\pi^{2})\over 768}+g^{5}\,{\pi(2\pi^{4}+5\pi^{2}-210)\over 15360}+g^{6}\,{\pi(105-\pi^{4})\over 10240}\dots$
@@
Figure 7: $A_{0}$ as a function of $g$ for the potential of squared Lorentz shape.
This concludes our presentation of results for the $s$-wave scattering length $A_{0}(g)$ for selected classes of short-range central potentials $V(r)\sim-g\,f(x)$. Of course, one could continue these analytical and numerical calculations for countless other profiles functions $f(x)$. A closed form expression for $A_{0}(g)$ can be expected to exist only in rare cases.

### v3

@@ Abstract
In these notes the Born series for the $s$-wave scattering $a_{0}$ is calculated for a class of central potentials $V(r)$ up to sixth order in a dimensionless coupling strength $g$. Examples of exponentially decaying potentials as well as truncated potentials involving a single length-scale $a$ are considered. In certain favorable cases the exact result for the $g$-dependent $s$-wave scattering length $a_{0}=A_{0}(g)a$ can be given in terms of special functions. The poles of $A_{0}(g)$ at increasing positive values of $g$ correspond to the thresholds, where $s$-wave bound-states occur successively. A scattering problem, where $A_{0}(g)$ is solvable in terms of elementary functions, is also presented.
@@ 2 Exponentially decaying potentials
In this section the Born series $A_{0}=\sum_{j=1}^{6}c_{j}g^{j}+\dots$ is computed for a variety of central potentials $V(r)$ with an exponential decay.
A: The familiar Yukawa potential is physically well motivated in terms of an underlying one-boson exchange mechanism. The pertinent profile function $f(x)$ and ”electrostatic potential” $\Phi(x)$ are in this case
@@
With this input one derives for the following Born series up to sixth order:
@@
The series for $A_{0}(g)$, with its first six terms written in eq.(8), has a finite radius of convergence $(|g|<g_{1})$, that is determined as that strength parameter $g_{1}$, where the threshold for the first $s$-wave bound-state in a Yukawa potential is reached. This particular value $g_{1}$ can be computed by solving numerically the second-order differential equation for the radial wave-function (multiplied with $x$):
@@
subject to a linear behavior $u(x)\sim x$ at small $x$. Note that $u^{\prime\prime}(x)+g\,x^{-1}u(x)=0$ has the regular solution in the form of a Bessel-function $u(x)=\sqrt{g\,x}J_{1}(2\sqrt{g\,x})=g\,x+{\cal O}(x^{2})$ [3], which also displays a linear behavior. Once the critical strength $g_{1}$ is reached, the asymptotic behavior of $u(x)$ for large $x$ turns into a constant. The convergence of this process is monitored by inspecting, how the first derivative $u^{\prime}(x)$ approaches zero for large $x$, when fine-tuning more and more digits of the strength parameter $g$. As a result of this numerical exercise, one finds that the first three thresholds $g_{1}$, $g_{2}$, $g_{3}$ for $s$-wave bound-states in the Yukawa potential lie at the values:
$g_{1}=1.679808\,,\qquad g_{2}=6.44726045\,,\qquad g_{3}=14.34202607\,.$ (12)
Note that the earlier determinations $g_{1}=1.680,\,g_{2}=6.445,\,g_{3}=14.34$ in ref.[4] are very close to the present results.
B: Next, we consider a purely exponentially decreasing potential $V(r)\sim-g\,e^{-r/a}$ with the profile function
@@
Figure 1: Scattering length $A_{0}=a_{0}/a$ versus $g$ for exponential potential $V(r)\sim-g\,e^{-r/a}$.
@@
Note that the numerical method, employed before to construct $g_{1},g_{2},g_{3}$ for the Yukawa potential, reproduces $g_{1}=1.4457965$, $g_{2}=7.6178156$, $g_{3}=18.7217517$ for the profile $f(x)=e^{-x}$.
C: As a further example, we choose an exponentially decreasing potential with a soft core:
@@
and derive the following Born series up to sixth order
@@
where all expansion coefficients are positive rational numbers. The radius of convergence $|g|<g_{1}$ of the series for $A_{0}(g)$ written in eq.(19) is determined by searching the threshold for the first $s$-wave bound-state in the potential $V(r)\sim-g\,x\,e^{-x}$. One finds with high numerical accuracy the values
$g_{1}=0.676681997\,,\qquad g_{2}=4.30765405\,,\qquad g_{3}=21.020136626\,.$ (20)
D: Next, we choose an exponentially decreasing potential with a very soft core, taking the profile function
@@
where all expansion coefficients are (large) rational numbers. The radius of convergence $|g|<g_{1}$ of the series for $A_{0}(g)$ in eq.(22) is again determined by searching the threshold for the first $s$-wave bound-state in the potential $V(r)\sim-g\,x^{2}e^{-x}$. One finds with high numerical accuracy the values
$g_{1}=0.216855259\,,\qquad g_{2}=1.580333075\,,\qquad g_{3}=8.027999134\,.$ (23)
E: Finally, one can consider an exponentially decreasing potential with a more singular behavior at small distances:
@@
The profile $f(x)=e^{-x}/x^{2}$ is too singular near the origin and therefore the coupling strength $g$ must be restricted to negative or small positive values $g<g_{\text{crit}}=1/4$. This feature concerning an instability is further elaborated on in the next section, when analyzing the truncated $1/r^{2}$-potential.
@@ 3 Truncated potentials
In this section we analyze short-range potentials, that vanish outside a sphere of radius $a$.
A: A well-known textbook problem [1] is the quantum mechanical scattering off a spherical potential-well, represented by the profile
@@
and now the knowledge of the ”electrostatic potential” $\Phi(x)$ is needed only in the inside region $0<x<1$. With little effort, one obtains the following Born series for the $s$-wave scattering length in a spherical potential-well up to sixth order:
@@
and, as a nice check, the Taylor expansion of the tangent-function confirms the series in eq.(27). The analytical result for $A_{0}(g)$ written in eq.(28) is shown in Fig. 2 in the region 40<g<62-40<g<62. The poles of $A_{0}(g)$ at the values $g_{j}=\pi^{2}(j-1/2)^{2}$ (where $\tan\sqrt{g_{j}}$ diverges) correspond to the well-known thresholds for the occurrence of successive $s$-wave bound-states in a spherical potential-well. The first six values of this infinite sequence read
@@
Figure 2: $A_{0}$ as a function of $g$ for spherical well potential.
B: It is also interesting to study the truncated $1/r^{2}$-potential with the profile function
@@
where the second term in curly brackets, $\sqrt{-g}J_{\nu+1}(\sqrt{-g})/J_{\nu}(\sqrt{-g})$, means $\sqrt{g}I_{\nu+1}(\sqrt{g})/I_{\nu}(\sqrt{g})$ for $0<g<1/4$. The exact result for $A_{0}(g)$ written in eq.(33) is shown in Fig. 3 for the parameter range 5<g<1/4-5<g<1/4. As already visible in Fig. 2, and valid for any truncated potential, the scattering length approaches with increasing repulsion the hard-core value: $\lim_{g\to-\infty}A_{0}(g)=-1$. On the positive side the boundary value, before instability sets in, is given by
@@
Figure 3: $A_{0}$ as a function of $g$ for truncated $1/r^{2}$-potential, restricted to range $g\leq 1/4$.
C: Next, we consider the truncated Coulomb-potential with profile function
@@
D: Let us form the difference of the truncated $1/r^{2}$-potential and the truncated Coulomb-potential by working with the profile function
$f(x)=\Big({1\over x^{2}}-{1\over x}\Big)\theta(1-x)\,,\qquad\Phi(x)={x\over 2}-\ln x\,.$ (39)
With this input one derives the following Born series up to sixth order
$A_{0}(g)={g\over 2}+{5\over 6}\,g^{2}+{275\over 144}\,g^{3}+{21949\over 4320}\,g^{4}+17.73814\,g^{5}\ +45.22973\,g^{6}+\dots\,$, (40)
where all expansion coefficients are positive rational numbers. The pertinent differential equation $u^{\prime\prime}(x)+g(x^{-2}-x^{-1})u(x)=0$ has the regular solution $u(x)=\sqrt{x}J_{\mu}(2\sqrt{-g\,x})$ involving a Bessel-function of index $\mu=\sqrt{1-4g}$. On this basis one obtains the following exact result for the $s$-wave scattering length
$A_{0}(g)=\bigg\{{1+\sqrt{1-4g}\over 2}-\sqrt{-g}\,{J_{\mu+1}(2\sqrt{-g})\over J_{\mu}(2\sqrt{-g})}\bigg\}^{-1}-1\,,\qquad\mu=\sqrt{1-4g}\,$, (41)
Figure 5: $A_{0}$ as a function of $g$ for difference of truncated $1/r^{2}$- and Coulomb-potential.
where the second term in curly brackets, $\sqrt{-g}J_{\mu+1}(2\sqrt{-g})/J_{\mu}(2\sqrt{-g})$, means $\sqrt{g}I_{\mu+1}(2\sqrt{g})/I_{\mu}(2\sqrt{g})$ for $0<g<1/4$. The expression for $A_{0}(g)$ written in eq.(41) is shown in Fig. 5 over the parameter range 5<g<1/4-5<g<1/4. The boundary value, where the slope diverges and instability sets in, is given by
$A_{0}(1/4)={I_{0}(1)-I_{1}(1)\over I_{0}(1)+I_{1}(1)}=0.38275296\,.$ (42)
E: A further interesting example is the truncated oscillator potential with the profile function
@@
Figure 6: $A_{0}$ as a function of $g$ for truncated oscillator potential.
F: As a further example we consider the truncated linear potential with the profile function
@@
Figure 7: $A_{0}$ as a function of $g$ for truncated linear potential.
G: As a last example we consider a truncated potential of cosinus shape with the profile function
$f(x)={\pi^{2}\over 4}\cos{\pi x\over 2}\,\theta(1-x)\,,\qquad\Phi(x)={\pi\over 2}+\cos{\pi x\over 2}-{4\over\pi x}\sin{\pi x\over 2}\,$, (50)
where the prefactor $\pi^{2}/4$ as been included for convenience. One derives the following Born series up to sixth order:
$A_{0}(g)=g\,{\pi^{2}-8\over 2\pi}+g^{2}\,{7\pi^{2}-66\over 24}+g^{3}\Big({\pi^{3}\over 6}-{17\pi\over 12}-{56\over 27\pi}\Big)+{g^{4}\over 39.594202}+{g^{5}\over 89.212563}+{g^{6}\over 201.00296}+\dots$ (51)
The differential equation $u^{\prime\prime}(x)+gf(x)u(x)=0$ for the radial wave-function is now solved by the odd Mathieu-function [3] with characteristic value 0 and parameter 2g-2g of the argument $\pi x/4$. The corresponding exact result for the $s$-wave scattering length $A_{0}(g)=u(1)/u^{\prime}(1)-1$ is shown in Fig. 8 over the parameter range 90<g<90-90<g<90. The first six poles, belonging to the thresholds for $s$-wave bound-states in the truncated cosinus-shaped potential, lie at values of the coupling strength:
$g_{1}=2.253065\,,\quad g_{2}=17.14564\,,\quad g_{3}=45.79166\,,\quad g_{4}=88.18843\,,\quad g_{5}=144.3357\,,\quad g_{6}=214.1899\,.$ (52)
Figure 8: $A_{0}$ as a function of $g$ for truncated potential of cosinus shape.
4 Further solvable potential scattering problems
A: Here, we present and discuss another solvable scattering problem, where the potential has the shape of a squared Lorentz-curve, i.e. the profile function is
@@ 4 Further solvable potential scattering problems
Figure 9: $A_{0}$ as a function of $g$ for potential of squared Lorentz shape.
@@
$A_{0}(g)$ $\!\!=\!\!$ $g\,{\pi\over 4}+g^{2}\,{\pi\over 16}+g^{3}\,{\pi(\pi^{2}-6)\over 192}+g^{4}\,{\pi(15-\pi^{2})\over 768}+g^{5}\,{\pi(2\pi^{4}+5\pi^{2}-210)\over 15360}+g^{6}\,{\pi(105-\pi^{4})\over 10240}\dots$ (57)
@@
B: Reconsider the $1/r^{2}$-potential, but now with a discontinuity at the boundary, by taking the profile function
$f(x)={\theta(1-x)\over x^{2}}\,,\qquad\Phi(x)=1-\ln x\,.$ (58)
The regular solution of $u^{\prime\prime}(x)+g\,x^{-2}u(x)=0$ is $u(x)=x^{(1+\sqrt{1-4g})/2}$ for restricted strength parameter $g\leq 1/4$, and this leads via $A_{0}=u(1)/u^{\prime}(1)-1$ to the following result for the $s$-wave scattering length
$A_{0}(g)={1-\sqrt{1-4g}\over 1+\sqrt{1-4g}}\,,\qquad\text{for}\qquad g\leq{1\over 4}\,.$ (59)
The corresponding Born series reads up to tenth order
$A_{0}(g)=g+2\,g^{2}+5\,g^{3}+14\,g^{4}+42\,g^{5}+132\,g^{6}+429\,g^{7}+1430\,g^{8}+4862\,g^{9}+16796\,g^{10}+\dots\$(60)
The expression for $A_{0}(g)$ written in eq.(59) is shown in Fig. 10 over the parameter range 6<g<1/4-6<g<1/4.
Figure 10: $A_{0}$ as a function of $g$ for $1/r^{2}$-potential with discontinuity at $r=a$.
C: In order that a truncated potential with a power-like drop-off $\sim r^{-\alpha}$ can support bound-states, the exponent $\alpha$ must be smaller than $2$. For this reason we consider the $n$-dependent profile function
$f(x)={\theta(1-x)\over x^{2-1/n}}\,,\qquad\Phi(x)=n-{n^{2}\over n+1}\,x^{1/n}\,.$ (61)
Figure 11: $A_{0}$ as a function of $g$ for profiles $f(x)=x^{-1}$ (dark) and $f(x)=x^{-3/2}$ (light).
The regular solution of $u^{\prime\prime}(x)+g\,x^{1/n-2}\,u(x)=0$ is given by a Bessel-function $u(x)=\sqrt{x}J_{n}\big(2n\sqrt{g\,x^{1/n}}\big)$, and this leads via $A_{0}=u(1)/u^{\prime}(1)-1$ to the following exact result for the $s$-wave scattering length
$A_{0}(g)={J_{n}(2n\sqrt{g})\over\sqrt{g}\,J_{n-1}(2n\sqrt{g})}-1\,,\qquad\text{for}\qquad g>0\,$,
$A_{0}(g)={I_{n}(2n\sqrt{-g})\over\sqrt{-g}\,I_{n-1}(2n\sqrt{-g})}-1\,,\qquad\text{for}\qquad g<0\,$, (62)
where $I_{n}(z)$ are modified Bessel-functions obtained from $J_{n}(z)$ by extension to imaginary argument. The curves in Figs. 11 and 12 display these results for $n=1,2$ and $n=3,4$, respectively.
Figure 12: $A_{0}$ as a function of $g$ for profiles $f(x)=x^{-5/3}$ (dark) and $f(x)=x^{-7/4}$ (light).
The Taylor expansion around $g=0$ gives the following Born series up to sixth order
$A_{0}(g)$ $=$ $g\,{n\over n\!+\!1}+g^{2}{2n^{2}\over(n\!+\!1)(n\!+\!2)}+g^{3}{n^{3}(5n\!+\!6)\over(n\!+\!1)^{2}(n\!+\!2)(n\!+\!3)}+g^{4}{2n^{4}(7n\!+\!12)\over(n\!+\!1)^{2}(n\!+\!2)(n\!+\!3)(n\!+\!4)}$
$+g^{5}{2n^{5}(21n^{3}\!+\!118n^{2}\!+\!214n\!+\!120)\over(n\!+\!1)^{3}(n\!+\!2)^{2}(n\!+\!3)(n\!+\!4)(n\!+\!5)}+g^{6}{4n^{6}(33n^{3}\!+\!230n^{2}\!+\!552n\!+\!360)\over(n\!+\!1)^{3}(n\!+\!2)^{2}(n\!+\!3)(n\!+\!4)(n\!+\!5)(n\!+\!6)}+\dots$
The poles of $A_{0}(g)$, which demarcate the thresholds for successive $s$-wave bound-states in a truncated potential with $r^{1/n-2}$ drop-off, occur at values of the coupling strength: $g_{j}^{(n)}=\big(z_{j}^{(n-1)}/2n\big)^{2}$, where $z_{j}^{(n-1)}$ denotes the $j$-th non-trivial zero of the Bessel-function $J_{n-1}(z)$. Interestingly, for $n=1$ these are the same values as found for the exponentially decreasing potential $f(x)=e^{-x}$ in sect. 2, eq.(17).
This concludes our presentation of results for the $s$-wave scattering length $A_{0}(g)$ for selected classes of short-range central potentials $V(r)\sim-g\,f(x)$. Of course, one could continue these analytical and numerical calculations for countless other profiles functions $f(x)$. A closed-form expression for $A_{0}(g)$ can be expected to exist only in rare cases.